Please show all the works and answer all the numbers accordingly (like the way the question is aking ). Thanks Check my work Suppose there are two independent economic factors, M, and M2. The risk-free rate is 4%, and all stocks have in
Answer #1

E(rp) = rf + (beta1 * RP1) + (beta2 * RP2)
32% = 4% + (1.7 * RP1) + (1.9 * RP2) => First equation
13% = 4% + (1.8 * RP1) + ((-0.7) * RP2) => second equation
Multiply first equation by 1.8 and second equation by 1.7
57.6% = 7.2% + (3.06 * RP1) + (3.42 *RP2)
22.1% = 6.8% + (3.06 * RP1) + (-1.19 * RP2)
Substract Equation 2 from equation 1
35.5% =0.4% + (4.61 * RP2)
RP2 = (35.5% - 0.4%) / 4.61 = 7.61%
RP1 = (32% - 4% - (1.9*7.61))/1.7 = 7.96

Expected return-beta relationship E(rp) = 4% + 7.96 BP1 + 7.61 BP2